You are the manager of Everyday Tomatoes; hence your firm operates in a perfectly competitive market....

  1. You are the manager of Everyday Tomatoes; hence your firm operates in a perfectly competitive market. The price in your market is $30 (per bushel). Your total cost curve is: C(Q) = 600 + 3Q2 (Q is 1 bushels).
    1. What level of output should you produce in the short run?
    2. What price should you charge in the short run?
    3. Will you make any profits in the short run?
    4. What will happen in the long run?
Answer #1

Given

C(Q)=600+3Q2

Marginal Cost is obtained by differentiating C(Q) with respect to Q, we get

MC(Q)=6Q

Market price=P=$30

a. What level of output should you produce in the short run?

A perfectly competitive firm sets its output level such that MC=Market price. So,

MC=P

6Q=30

Q=30/6=5 bushel

Level of output in short run=5 bushel

b. What price should you charge in the short run?

Competitive firm is price taker. It will charge the same price as Market price i..e $30 per bushel

c. Will you make any profits in the short run?

Total Revenue=TR=P*Q=30*5=$150

Total Cost=600+3Q2=600+3*52=$675

Profit=TR-TC=150-675=-$525

Firm will make a loss of $525

d. What will happen in the long run?

Firm is making a loss. Firms will start exiting the market. This will start decreasing supply and price will start increasing. In long run firm will operate at minimum ATC and profit will be zero in long run.